OpenStudy (anonymous):

I am actually bored, lets try solving this one: Here is a function \(f\) such that \(f(2) = 60 \) and \( −f(1) + f(2) − f(3) + f(4)−f(5) \cdots +f(n) = nf(n)\), for all positive integers \( n > 1 \) Can you find \( f(11) \) ?

OpenStudy (anonymous):

Is this too hard ?

OpenStudy (mr.math):

No. Give us some time.

OpenStudy (anonymous):

Sure :)

OpenStudy (mr.math):

\[f(2)-f(11)=11f(11) \implies 60=12f(11) \implies f(11)=5.\]

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

Good question. This one takes a while due to its recursive nature ...

OpenStudy (anonymous):

Thanks, one could try solving it by general analysis and breaking it into smaller cases.

OpenStudy (anonymous):

ksaimouli, please don't post your questions in other questions. Yeah I give up FoolForMath :-( Do you have an answer?

OpenStudy (anonymous):

Yes, 10.

OpenStudy (anonymous):

Haha what a stupid question. "Does FFM have an answer to a mathematical problem?" is the constant value of "YES" :-P

OpenStudy (anonymous):

lol, I am not that smart buddy, check out my username -> FOOOOOL :D

OpenStudy (anonymous):

I don't believe THAT for a single second! ;)

OpenStudy (anonymous):

Yes, believe it after that single second past :P

OpenStudy (anonymous):

LIES!

OpenStudy (mr.math):

Wait, where did I go wrong. We can easily show that f(1)=0, and that f(2n-1)=f(2n) for \(n\ge 2\). From here we can reduce \(-f(1)+f(2)-f(3)+..-f(11)=f(2)-f(11)\). Doesn't this sound right?

OpenStudy (anonymous):

This is a GRE problem, and \( 5 \) is not even in the four given option. Something wrong in you analysis I believe, I got different f(2k) and f(2k+1) for n=2k and n=2k+1 respectively.

OpenStudy (mr.math):

Could you give a counterexample?

OpenStudy (mr.math):

Oh the definition says for all n>1.

OpenStudy (mr.math):

Isn't f(1) given by the definition?

OpenStudy (nikvist):